OpenStudy (anonymous):

Solve for x and y: 3/x=2/5 y/4=y-1/3

OpenStudy (anonymous):

solve for x on top one then plug it into bottom one.

OpenStudy (anonymous):

Or you can use matrices. Set up augmented matrix and reduce to echelon form.

OpenStudy (anonymous):

yea...how do u do that? lol

OpenStudy (asnaseer):

top equation is:\[\frac{3}{x}=\frac{2}{5}\]just cross multiply to find x.

OpenStudy (asnaseer):

is the bottom equation this?\[\frac{y}{4}=y-\frac{1}{3}\]

OpenStudy (anonymous):

no its y-1 on top and 3 on bottom

OpenStudy (asnaseer):

so botton equation is:\[\frac{y}{4}=\frac{y-1}{3}\]again, just cross multiply and solve to find y

OpenStudy (asnaseer):

by cross multiply I mean this: lets say you were given:\[\frac{a}{b}=\frac{c}{d}\]then you can say:\[a\times d=c\times b\]

OpenStudy (anonymous):

so for x it would be 15 over 2 which equals 7.5?

OpenStudy (asnaseer):

It's called cross multiplication because of this diagram: